Please login if you have access to particular applications.
|Subject:||ADCP external trigger restarted|
|When:||2015-12-08 05:26:28 UTC [entered into Elog as 2015-12-08 05:26:28]|
Position from underway logging
|Equipment:||ADCP Linked to equipment register: 75 KHz / 150 KHz ADCP [view]|
GSM noted that the external triggering for the ADCP had been switched off - it has now been restarted. It is unknown how long triggering as been off.
|Logged by:||Steven Van Graas [entered into Elog as Steve van Grass]|